I need to find det (lambda * x - A): The determinant for this answer is: -x^3+7x-6. How can I get this answer?8. (10 points) Find the general solution to the homogeneous system of DE: 3 2 6 x = Ax where A = -2 -1 -2 -4) 1 -2
Answer #1

a) Given system, X= AX Where Aa 2 2. nl .22 Let x be eigenvalue of A. Then det (A-XI) = 0 a ♡ 3-1 2 I-a. -2 6 2. 4- 2 x3+72-so АУ - 2y. а. - 2 -2 22 - ) М4 а, 5 - 2 - 2 -2 -V - No- 6 This gives N +4, 16V = 6 -2 - 2 - 3 = -NA-V2 - 6N3 solving this g